### v4

# Hamiltonian stationary Lagrangian surfaces with harmonic mean curvature in complex space forms

Abstract.
In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms whose mean curvature is a harmonic function. We prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel. In the non-constant mean curvature case, under the condition of constant Gaussian curvature, we give an explicit description of all such surfaces.
Key words and phrases:
Hamiltonian stationary Lagrangian surfaces, complex space forms, harmonic mean curvature
2010 Mathematics Subject Classification:
Primary: 53C42; Secondary: 53B25
1. Introduction
[…]
(1.1) ${\rm div}(JH)=0$
on $M$ (cf. [9]), where div is the divergence on $M$.
In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms. First, we prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel (Proposition 3.1). Next, we investigate the case where the mean curvature is a non-constant harmonic function. Our purpose is to give an explicit description of all such surfaces under the condition of constant Gaussian curvature (Theorem 4.1).
2. Preliminaries
Let $\tilde{M}^{n}(4\epsilon)$ be a complete and simply connected complex space form of complex dimension $n$ and constant holomorphic sectional curvature $4\epsilon$, that is, $\tilde{M}^{n}(4\epsilon)$ is the complex Euclidean space $\mathbb{C}^{n}$, the complex projective space $\mathbb{C}P^{n}(4\epsilon)$ or the complex hyperbolic space $\mathbb{C}H^{n}(4\epsilon)$ according as $\epsilon=0$, $\epsilon>0$ or $\epsilon<0$.
Let $M$ be a Lagrangian submanifold of $\tilde{M}^{n}(4\epsilon)$. We denote the Levi-Civita connections on $M^{n}$ and $\tilde{M}^{n}(4\epsilon)$ by $\nabla$ and $\tilde{\nabla}$, respectively. The Gauss and Weingarten formulas are respectively given by
$\tilde{\nabla}_{X}Y=\nabla_{X}Y+h(X,Y),\quad\tilde{\nabla}_{X}\xi=-A_{\xi}X+D_{X}\xi$
for tangent vector fields $X$, $Y$ and normal vector field $\xi$, where $h,A$ and $D$ are the second fundamental form, the shape operator and the normal connection. The mean curvature vector field $H$ is defined by $H=(1/n){\rm trace}\,h.$ The function $|H|$ is called the mean curvature. We have (cf. [7])
(2.1) $D_{X}JY=J(\nabla_{X}Y),$
(2.2) $\left<h(X,Y),JZ\right>=\left<h(Y,Z),JX\right>=\left<h(Z,X),JY\right$>.
Denote by $R$ the Riemann curvature tensor of $\nabla$. Then the equations of Gauss and Codazzi are respectively given by
$\left<R(X,Y)Z,W\right$>= $\left<h(Y,Z),h(X,W)\right>-\left<h(X,Z),h(Y,W)\right$>
[…]
(2.4) $({\bar{\nabla}}_{X}h)(Y,Z)=$ $({\bar{\nabla}}_{Y}h)(X,Z),$
where $X,Y,Z,W$ are vectors tangent to $M$, and $\bar{\nabla}h$ is defined by
$({\bar{\nabla}}_{X}h)(Y,Z)=D_{X}h(Y,Z)-h(\nabla_{X}Y,Z)-h(Y,\nabla_{X}Z).$
3. Hamiltonian stationary Lagrangian surfaces
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H\neq 0$ everywhere. Denote by $K$ the Gaussian curvature of $M$. Let $\{e_{1},e_{2}\}$ be a local orthonormal basis of $M$ such that $Je_{1}$ is parallel to $H$. It follows from (2.2) that second fundamental form takes the form
(3.1) $\begin{split}&h(e_{1},e_{1})=(a-c)Je_{1}+bJe_{2},\\ &h(e_{1},e_{2})=bJe_{1}+cJe_{2},\\ &h(e_{2},e_{2})=cJe_{1}-bJe_{2}\end{split}$
for some functions $a$, $b$ and $c$. Putting $\omega_{i}^{j}(e_{k})=\left<\nabla_{e_{k}}e_{i},e_{j}\right$>, by (2.1) and (LABEL:sf) we have
$(\bar{\nabla}_{e_{1}}h)(e_{2},e_{2})$ $=\{e_{1}c+3b\omega_{1}^{2}(e_{1})\}Je_{1}-\{e_{1}b-3c\omega_{1}^{2}(e_{1})\}Je_{2},$
$(\bar{\nabla}_{e_{2}}h)(e_{1},e_{2})$ $=\{e_{2}b+(a-3c)\omega_{1}^{2}(e_{2})\}Je_{1}+\{e_{2}c+3b\omega_{1}^{2}(e_{2})\}Je_{2},$
$(\bar{\nabla}_{e_{1}}h)(e_{1},e_{2})$ $=\{e_{1}b+(a-3c)\omega_{1}^{2}(e_{1})\}Je_{1}+\{e_{1}c+3b\omega_{1}^{2}(e_{1})\}Je_{2},$
$(\bar{\nabla}_{e_{2}}h)(e_{1},e_{1})$ $=\{e_{2}(a-c)-3b\omega_{1}^{2}(e_{2})\}Je_{1}+\{e_{2}b+(a-3c)\omega_{1}^{2}(e_{2})\}Je_{2}.$
Therefore, the equation (2.4) of Codazzi implies
(3.2) $e_{1}c+3b\omega_{1}^{2}(e_{1})=e_{2}b+(a-3c)\omega^{2}_{1}(e_{2}),$
(3.3) $-e_{1}b+3c\omega_{1}^{2}(e_{1})=e_{2}c+3b\omega^{2}_{1}(e_{2}),$
(3.4) $e_{2}(a-c)-3b\omega_{1}^{2}(e_{2})=e_{1}b+(a-3c)\omega^{2}_{1}(e_{1}).$
Combining (3.3) and (3.4) yields
(3.5) $e_{2}a-a\omega^{2}_{1}(e_{1})=0.$
The Hamiltonian stationary condition (1.1) is equivalent to
(3.6) $e_{1}a+a\omega_{1}^{2}(e_{2})=0.$
Using (3.5) and (3.6), we obtain
$a^{-1}e_{1},a^{-1}e_{2}]=0.$
Therefore, there exists a local coordinate system $\{u,v\}$ such that $e_{1}=a\partial_{u}$, $e_{2}=a\partial_{v}$. Hence, the metric tensor is given by
(3.7) $g=a^{-2}(du^{2}+dv^{2}),$
which implies that
(3.8) $\omega_{1}^{2}(e_{1})=a_{v},\quad\omega_{1}^{2}(e_{2})=-a_{u},$
(3.9) $K=-(a_{u})^{2}-(a_{v})^{2}+a(a_{uu}+a_{vv}).$
Equations (3.2) and (3.3) are rewritten as
(3.10) $ac_{u}+3ba_{v}$ $=ab_{v}-(a-3c)a_{u},$
(3.11) $-ab_{u}+3ca_{v}$ $=ac_{v}-3ba_{u}.$
Put $\delta=\epsilon-K$. Then the equation (2.3) of Gauss and (LABEL:sf) yield that
(3.12) $\delta=2b^{2}-ac+2c^{2}.$
Proposition 3.1.
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$. If $|H|$ is non-zero constant, then $\bar{\nabla}h=0$.
Proof. From (3.5) and (3.6), it follows that if $|H|=|a|/2$ is non-zero constant, then $\omega_{1}^{2}=0$. Hence we have $K=0$. Differentiating (3.12) along $e_{i}$ $(i=1,2)$ yields $4be_{i}b+(4c-a)e_{i}c=0$, which together with (3.2) and (3.3), shows that $b$ and $c$ are constant. Thus, using $\omega_{1}^{2}=0$, (2.1) and (LABEL:sf), we obtain $\bar{\nabla}h=0$.
Remark 3.1.
An explicit description of all Lagrangian surfaces with $\bar{\nabla}h=0$ in $\tilde{M}^{2}(4\epsilon)$ has been obtained in Theorem 7.2 in [1], Theorem A.1 and Theorem A.2 in [4].
4. Main theorem
Let ${\mathbb{C}}^{3}_{1}$ be the complex 33-space endowed with the inner product
$\left<(z_{1},z_{2},z_{3}),(w_{1},w_{2},w_{3})\right>={\rm Re}(-z_{1}\bar{w}_{1}+z_{2}\bar{w}_{2}+z_{3}\bar{w}_{3}).$
Put $H_{1}^{5}(-1)=\{z\in{\mathbb{C}}^{3}_{1}:\left<z,z\right>=-1\}$. The map $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ defined by $z\mapsto z\cdot{\mathbb{C}}^{*}$ is referred to as the Hopf fibration.
The main result of this paper is the following classification theorem.
Theorem 4.1.
Let $M$ be a Hamiltonian stationary Lagrangian surface in a 22-dimensional complex space form $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H$ is nowhere vanishing. If $M$ has non-constant harmonic mean curvature and constant Gaussian curvature $K$, then $K=\epsilon=-1$ and $M$ is locally congruent to the image of $\Pi\circ\phi$, where $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ is the Hopf fibration and $\phi:M\rightarrow H_{1}^{5}(-1)\subset{\mathbb{C}}^{3}_{1}$ is given by one of the following immersions::
$(1)$
$\phi(x,y)=\biggl(me^{y}+\dfrac{e^{-y}+2im^{2}xe^{y}}{2m},me^{ix+y},\dfrac{e^{-y}+2im^{2}xe^{y}}{2m}\biggr);$
$(2)$
$\phi(x,y)=\biggl(1-\dfrac{i(1+m^{2})}{m^{2}x+y},\dfrac{m\sqrt{1+m^{2}}e^{ix}}{m^{2}x+y},\dfrac{\sqrt{1+m^{2}}e^{iy}}{m^{2}x+y}\biggr),$
where $m$ is a positive real number.
Proof. We shall use the same notation as in Section 3. Suppose that $|H|$ is a harmonic function on $M$. Then by (3.7) we have
(4.1) $a_{uu}+a_{vv}=0.$
Moreover, suppose that $|H|$ is non-constant and $K$ is constant. Then, combining (3.9) and (4.1) shows that $K<0$ and
(4.2) $a_{u}=\sqrt{-K}\cos\theta,\quad a_{v}=\sqrt{-K}\sin\theta.$
for some function $\theta$ on $M$. Substituting (4.2) into (4.1), we have
(4.3) $(\sin\theta)\theta_{u}+(\cos\theta)\theta_{v}=0.$
On the other hand, since $a_{uv}-a_{vu}=0$ holds, by (4.2) we obtain
(4.4) $(\cos\theta)\theta_{u}+(\sin\theta)\theta_{v}=0.$
It follows (4.3) from (4.4) that $\theta_{u}=\theta_{v}=0$, that is, $\theta$ is constant. Solving (4.2), we conclude that up to translations, $a$ is given by
(4.5) $a=\sqrt{-K}\{(\cos\theta)u+(\sin\theta)v\}.$
Case (i): $b=0$ on an open subset $\mathcal{U}$. In this case, (3.10) and (3.11) reduce respectively to
(4.6) $(a-3c)a_{u}+ac_{u}=0,$
(4.7) $3ca_{v}-ac_{v}=0.$
Differentiating (3.12) with respect to $v$, we have
(4.8) $ca_{v}+(a-4c)c_{v}=0.$
Combining (4.7) and (4.8) gives
$(a-c)c_{v}=0.$
If $c_{v}\neq 0$ on an open subset in $\mathcal{U}$, then $a=c$ on that open subset. From (3.12) we see that $a$ is constant, which contradicts our assumption. Hence we have $c_{v}=0$ on $\mathcal{U}$. Thus, by (4.7) we get
$ca_{v}=0.$
Case (i.1): $c=0$ on an open subset $\mathcal{U}_{1}\subset\mathcal{U}$. In this case, from (3.12) we have $K=\epsilon=-1$. It follows from (4.6) that $a_{u}=0$ on $\mathcal{U}_{1}$. Hence, using (4.5) and K=1K=-1 yields that $a^{2}=v^{2}$. Applying the coordinate transformation: $y=-\int v^{-1}dv$, we see that the metric tensor (3.7) becomes
$g=m^{2}e^{2y}du^{2}+dy^{2}$
for some positive constant $m$, and the second fundamental form satisfies
$h(\partial_{u},\partial_{u})=J\partial_{u},\quad h(\partial_{u},\partial_{y})=h(\partial_{y},\partial_{y})=0.$
Rewriting $u$ with $x$ and according to [3, page 3475], we conclude that $\mathcal{U}_{1}$ is congruent to the Lagrangian surface obtained from (1).
Case (i.2): $a_{v}=0$ on an open subset $\mathcal{U}_{2}\subset\mathcal{U}$. In this case, $a_{u}\neq 0$. Differentiating (3.12) with respect to $u$ leads to
(4.9) $ca_{u}+(a-4c)c_{u}=0.$
Eliminating $c_{u}$ from (4.6) and (4.9) gives
$(a-2c)(a-6c)a_{u}=0.$
If $a\neq 2c$ on an open set in $\mathcal{U}_{2}$, then $a=6c$ and it follows from (3.12) that $a$ is constant, which contradicts our assumption. Hence $a=2c$, which together with (3.12) implies $K=\epsilon=-1$. Thus, by (4.5) we obtain $a^{2}=u^{2}$. Applying the coordinate transformation: $x=(u+v)/2$ and y=(uv)/2y=(u-v)/2, we see that the metric tensor (3.7) becomes
$g=\dfrac{2}{(x+y)^{2}}(dx^{2}+dy^{2})$
and the second fundamental form satisfies
$h(\partial_{x},\partial_{x})=J\partial_{x},\quad h(\partial_{x},\partial_{y})=0,\quad h(\partial_{y},\partial_{y})=J\partial_{y}.$
According to [5, page 124], $\mathcal{U}_{2}$ is congruent to the Lagrangian surface obtained from (2) with $m=1$.
Case (ii): $b\neq 0$ on an open subset $\mathcal{V}$. We put $A=a_{u}$ and $B=a_{v}$, which are constants. Differentiating (3.12) with respect to $u$ and $v$, we obtain
(4.10) $b_{u}=\frac{cA+(a-4c)c_{u}}{4b},$
(4.11) $b_{v}=\frac{cB+(a-4c)c_{v}}{4b}.$
Substituting (4.10) and (4.11) into (3.10) and (3.11) yields
(4.12) $4abc_{u}-a(a-4c)c_{v}=4b(3c-a)A-(12b^{2}-ac)B,$
(4.13) $a(a-4c)c_{u}+4abc_{v}=(12b^{2}-ac)A+12bcB.$
Solving (4.12) and (4.13) for $c_{u}$ and $c_{v}$ and using (3.12), we have
(4.14) $c_{u}=\frac{Af_{1}+Bf_{2}}{a(a^{2}+8\delta)},$
(4.15) $c_{v}=\frac{Af_{3}+Bf_{4}}{a(a^{2}+8\delta)},$
where
$f_{1}=-4ab^{2}-a^{2}c+4ac^{2},$
$f_{2}=-48b^{3}+16abc-48bc^{2},$
$f_{3}=4a^{2}b+48b^{3}-32abc+48bc^{2},$
$f_{4}=12ab^{2}-a^{2}c+4ac^{2}.$
Substituting (4.14) and (4.15) into (4.10) and (4.11) gives
(4.16) $b_{u}=$ $\frac{Ag_{1}+Bg_{2}}{a(a^{2}+8\delta)},$
(4.17) $b_{v}=$ $\frac{Ag_{3}+Bg_{4}}{a(a^{2}+8\delta)},$
where
$g_{1}=-a^{2}b+8abc,$
$g_{2}=-12ab^{2}+4a^{2}c+48b^{2}c-28ac^{2}+48c^{3},$
$g_{3}=a^{3}+12ab^{2}-12a^{2}c-48b^{2}c+44ac^{2}-48c^{3},$
$g_{4}=3a^{2}b-8abc.$
Differentiating (4.14) and (4.15) with respect to $v$ and $u$ respectively, we have
(4.18) $c_{uv}=\frac{(Af_{1}+Bf_{2})_{v}a(a^{2}+8\delta)-(Af_{1}+Bf_{2})(3a^{2}+8\delta)B}{a^{2}(a^{2}+8\delta)^{2}},$
(4.19) $c_{vu}=\frac{(Af_{3}+Bf_{4})_{u}a(a^{2}+8\delta)-(Af_{3}+Bf_{4})(3a^{2}+8\delta)A}{a^{2}(a^{2}+8\delta)^{2}}.$
By a straightforward but tedious computation, using (4.14)-(4.17) and substituting (3.12) into the numerators of (4.18) and (4.19), we get
(4.20) $c_{uv}=\frac{A^{2}h_{1}+ABh_{2}+B^{2}h_{3}}{a^{2}(a^{2}+8\delta)^{2}},$
(4.21) $c_{vu}=\frac{A^{2}h_{4}+ABh_{5}+B^{2}h_{6}}{a^{2}(a^{2}+8\delta)^{2}},$
where
$h_{1}=$ $-12a^{4}b-144a^{2}b^{3}+160a^{3}bc+768ab^{3}c-656a^{2}bc^{2}+768abc^{3},$
$h_{2}=$ $-108a^{3}b^{2}-960ab^{4}+18a^{4}c+1168a^{2}b^{2}c+2304b^{4}c-252a^{3}c^{2}$
$-3840ab^{2}c^{2}+1296a^{2}c^{3}+4608b^{2}c^{3}-2880ac^{4}+2304c^{5},$
$h_{3}=$ $-96a^{2}b^{3}+768b^{5}-384ab^{3}c+32a^{2}bc^{2}+1536b^{3}c^{2}-384abc^{3}+768bc^{4},$
$h_{4}=$ $-8a^{4}b-96a^{2}b^{3}-768b^{5}+128a^{3}bc+1152ab^{3}c-608a^{2}bc^{2}$
$-1536b^{3}c^{2}+1152abc^{3}-768bc^{4},$
$h_{5}=$ $-92a^{3}b^{2}-192ab^{4}+18a^{4}c+784a^{2}b^{2}c+2304b^{4}c-252a^{3}c^{2}$
$-3072ab^{2}c^{2}+1296a^{2}c^{3}+4608b^{2}c^{3}-2880ac^{4}+2304c^{5},$
$h_{6}=$ $-240a^{2}b^{3}+80a^{3}bc+768ab^{3}c-496a^{2}bc^{2}+768abc^{3}.$
Since $c_{uv}-c_{vu}$=0 holds, from (4.20), (4.21), it follows that
(4.22) $4b(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})=0,$
where
$k_{1}=$ $-a^{4}-12a^{2}b^{2}+192b^{4}+8a^{3}c-96ab^{2}c-12a^{2}c^{2}$
$+384b^{2}c^{2}-96ac^{3}+192c^{4},$
$k_{2}=$ $-4a^{3}b-192ab^{3}+96a^{2}bc-192abc^{2},$
$k_{3}=$ $36a^{2}b^{2}+192b^{4}-20a^{3}c-288ab^{2}c$
$+132a^{2}c^{2}+384b^{2}c^{2}-288ac^{3}+192c^{4}.$
By a straightforward but tedious computation, using (4.14)-(4.17), we get
(4.23) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{u}$ $=\frac{4(A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4})}{a^{2}+8\delta},$
(4.24) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{v}$ $=\frac{8(A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8})}{a^{2}+8\delta},$
where
$P_{1}=$ $-a^{5}-24a^{3}b^{2}-192ab^{4}+12a^{4}c+168a^{2}b^{2}c+384b^{4}c$
$-56a^{3}c^{2}-576ab^{2}c^{2}+168a^{2}c^{3}+768b^{2}c^{3}-384ac^{4}+384c^{5},$
$P_{2}=$ $-2a^{4}b-72a^{2}b^{3}-1920b^{5}+24a^{3}bc+1344ab^{3}c$
$-264a^{2}bc^{2}-3840b^{3}c^{2}+1344abc^{3}-1920bc^{4},$
$P_{3}=$ $32a^{3}b^{2}+960ab^{4}-14a^{4}c-888a^{2}b^{2}c-2688b^{4}c+224a^{3}c^{2}$
$+4032ab^{2}c^{2}-1272a^{2}c^{3}-5376b^{2}c^{3}+3072ac^{4}-2688c^{5},$
$P_{4}=$ $24a^{2}b^{3}+1152b^{5}-8a^{3}bc-960ab^{3}c$
$+216a^{2}bc^{2}+2304b^{3}c^{2}-960abc^{3}+1152bc^{4},$
$P_{5}=$ $a^{4}b+60a^{2}b^{3}+576b^{5}-32a^{3}bc-672ab^{3}c$
$+252a^{2}bc^{2}+1152b^{3}c^{2}-672abc^{3}+576bc^{4},$
$P_{6}=$ $-a^{5}-38a^{3}b^{2}-192ab^{4}+24a^{4}c+444a^{2}b^{2}c+1344b^{4}c-218a^{3}c^{2}$
$-2016ab^{2}c^{2}+924a^{2}c^{3}+2688b^{2}c^{3}-1824ac^{4}+1344c^{5},$
$P_{7}=$ $-4a^{4}b-180a^{2}b^{3}-960b^{5}+96a^{3}bc+1248ab^{3}c$
$-564a^{2}bc^{2}-1920b^{3}c^{2}+1248abc^{3}-960bc^{4},$
$P_{8}=$ $6a^{3}b^{2}-5a^{4}c-84a^{2}b^{2}c-192b^{4}c+50a^{3}c^{2}$
$+288ab^{2}c^{2}-180a^{2}c^{3}-384b^{2}c^{3}+288ac^{4}-192c^{5}.$
Thus, it follows from (4.22), (4.23) and (4.24) that
(4.25) $A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4}=0,$
(4.26) $A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8}=0.$
Using a computer algebra system, we check that the resultant of the left-hand sides of (4.25) and (4.26) with respect to AA is, up to a constant factor,
(4.27) $B^{9}(2b^{2}-ac+2c^{2})^{3}(a^{2}+16b^{2}-8ac+16c^{2})^{7}(a^{2}+48b^{2}-24ac+48c^{2})^{2}Q,$
where
$Q=$ $5a^{6}+120a^{4}b^{2}+720a^{2}b^{4}+2304b^{6}-60a^{5}c-720a^{3}b^{2}c$
$-3456ab^{4}c+300a^{4}c^{2}+3168a^{2}b^{2}c^{2}+6912b^{4}c^{2}-1008a^{3}c^{3}$
$-6912ab^{2}c^{3}+2448a^{2}c^{4}+6912b^{2}c^{4}-3456ac^{5}+2304c^{6}.$
Combining (3.12) and (4.27), we find that (4.27) can be simplified to
(4.28) $B^{9}\delta^{3}(a^{2}+8\delta)^{7}(a^{2}+24\delta)^{2}(5a^{6}+60a^{4}\delta+180a^{2}\delta^{2}+288\delta^{3}),$
which must vanish on $\mathcal{V}$. Taking into account $\nabla a\neq 0$, we have
$B\delta=0.$
If $\delta\neq 0$, then $B=0$, which implies that (4.22) and (4.26) reduce respectively to
$a^{4}+12a^{2}b^{2}-192b^{4}-8a^{3}c+96ab^{2}c+12a^{2}c^{2}$
(4.29) $-384b^{2}c^{2}+96ac^{3}-192c^{4}=0,$
$a^{4}+60a^{2}b^{2}+576b^{4}-32a^{3}c-672ab^{2}c$
(4.30) $+252a^{2}c^{2}+1152b^{2}c^{2}-672ac^{3}+576c^{4}=0.$
Subtracting (4.29) from (4.30), we derive
(4.31) $24(2b^{2}-ac+2c^{2})(a^{2}+16b^{2}-8ac+16c^{2})=0.$
Combining (3.12) and (4.31) implies
$a^{2}+8\delta=0,$
which contradicts $\nabla a\neq 0$. Hence $\delta=0$, that is, $K=\epsilon=-1$.
From (3.12) we see that $c\neq 0$. Changing the sign of $e_{1}$ if necessary, we may assume that $c>0$. Put
(4.32) $f=\frac{\sqrt{b^{2}+c^{2}-b\sqrt{b^{2}+c^{2}}}}{\sqrt{2}(b^{2}+c^{2})},\quad k=\frac{\sqrt{b^{2}+c^{2}+b\sqrt{b^{2}+c^{2}}}}{\sqrt{2}(b^{2}+c^{2})},$
which are non-zero and unequal everywhere. It follows from (3.12) and (4.32) that
(4.33) $a=\frac{\sqrt{f^{2}+k^{2}}}{fk},\quad b=\frac{k^{2}-f^{2}}{(f^{2}+k^{2})^{\frac{3}{2}}},\quad c=\frac{2fk}{(f^{2}+k^{2})^{\frac{3}{2}}}.$
We make the following change of the basis:
(4.34) $\tilde{e}_{1}=\frac{k}{\sqrt{f^{2}+k^{2}}}e_{1}+\frac{f}{\sqrt{f^{2}+k^{2}}}e_{2},\quad\tilde{e}_{2}=\frac{f}{\sqrt{f^{2}+k^{2}}}e_{1}-\frac{k}{\sqrt{f^{2}+k^{2}}}e_{2}.$
Then, using (LABEL:sf), (4.33) and (4.34), we find
$h(\tilde{e}_{1},\tilde{e}_{1})=f^{-1}J\tilde{e}_{1},\quad h(\tilde{e}_{1},\tilde{e}_{2})=0,\quad h(\tilde{e}_{2},\tilde{e}_{2})=k^{-1}J\tilde{e}_{2}.$
Therefore, according to [5, Theorem 4.1], there exists a local coordinate system $\{x,y\}$ such that $\partial_{x}=f\tilde{e}_{1}$ and $\partial_{y}=k\tilde{e}_{2}$, where $f$ and $k$ satisfy
(4.35) $\frac{f_{y}}{k}=\frac{k_{x}}{f},\quad\Bigl(\frac{f_{y}}{k}\Bigr)_{y}+\Bigl(\frac{k_{x}}{f}\Bigr)_{x}=fk.$
The Hamiltonian stationary condition (1.1) is equivalent to (cf. [8])
(4.36) $\Bigl(\frac{k}{f}\Bigr)_{x}+\Bigl(\frac{f}{k}\Bigr)_{y}=0.$
By Theorem 3.1 of [2], up to translations and sign, the exact solutions of the over-determined PDE system (4.35)-(4.36) are given by
(4.37) $f=\lambda m\,{\rm csch}\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr),\quad$ $k=\lambda\,{\rm csch}\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr);$
(4.38) $f=\lambda m\,\sec\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr),\quad$ $k=\lambda\,\sec\Bigl(\frac{\lambda(m^{2}x+y)}{\sqrt{1+m^{2}}}\Bigr);$
(4.39) $f=\frac{m\sqrt{1+m^{2}}}{m^{2}x+y},\quad$ $k=\frac{\sqrt{1+m^{2}}}{m^{2}x+y},$
where $\lambda$ and $m$ are positive real numbers.
We note that $f=mk$. By the first equation of (4.33) we have
$a=\frac{\sqrt{1+m^{2}}}{mk}.$
Furthermore, it follows from (4.34) that
$\partial_{u}=\frac{1}{1+m^{2}}(\partial_{x}+m^{2}\partial_{y}),\quad\partial_{v}=\frac{m}{1+m^{2}}(\partial_{x}-\partial_{y}).$
Since $a_{u}$ and $a_{v}$ are constant, the solutions (4.37) and (4.38) are excluded. Considering also that $a_{u}^{2}+a_{v}^{2}=1$ and $f\neq k$, we see that $f$ and $g$ are given by (4.39) with $m\neq 1$. In [6, Section 7], it is proved that the corresponding surface is congruent to the Lagrangian surface obtained from case (2) with $m\neq 1$. This completes the proof.

### v5

@@
 Abstract.
-In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms whose mean curvature is a harmonic function.
-We prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel.
-In the non-constant mean curvature case, under the condition of constant Gaussian curvature, we give an explicit description of all such surfaces.
+In this paper, we study Hamiltonian stationary Lagrangian surfaces in complex space forms.
+We first show that when the mean curvature is a non-zero constant, the second fundamental form is parallel.
+We then consider the case in which the mean curvature is a non-constant harmonic function.
+Under the additional assumption that the Gaussian curvature is constant, we obtain a complete classification of such Lagrangian surfaces.
 Key words and phrases:
@@
 […]
-(1.1) ${\rm div}(JH)=0$
-on $M$ (cf. [9]), where div is the divergence on $M$.
-In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms.
-First, we prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel (Proposition 3.1).
-Next, we investigate the case where the mean curvature is a non-constant harmonic function.
-Our purpose is to give an explicit description of all such surfaces under the condition of constant Gaussian curvature (Theorem 4.1).
+(1.1) ${\rm div}(JH)=0,$
+where $H$ denotes the mean curvature vector field of $M$.
+In this paper, we investigate Hamiltonian stationary Lagrangian surfaces in complex space forms.
+When the mean curvature is a non-zero constant, we show that the second fundamental form is parallel.
+Such Lagrangian surfaces have been completely classified in Theorem 7.2 of [1], Theorems A.1 and A.2 of [4].
+We then turn to the case in which the mean curvature is a non-constant harmonic function.
+Assuming that the Gaussian curvature is constant, we prove that the ambient space must be the complex hyperbolic plane and the Gaussian curvature necessarily takes a specific negative value.
+This phenomenon leads to a complete classification.
 2. Preliminaries
@@
 We denote the Levi-Civita connections on $M^{n}$ and $\tilde{M}^{n}(4\epsilon)$ by $\nabla$ and $\tilde{\nabla}$, respectively.
-The Gauss and Weingarten formulas are respectively given by
-$\tilde{\nabla}_{X}Y=\nabla_{X}Y+h(X,Y),\quad\tilde{\nabla}_{X}\xi=-A_{\xi}X+D_{X}\xi$
-for tangent vector fields $X$, $Y$ and normal vector field $\xi$, where $h,A$ and $D$ are the second fundamental form, the shape operator and the normal connection.
+The Gauss and Weingarten formulas are given respectively by
+$\tilde{\nabla}_{X}Y=\nabla_{X}Y+h(X,Y),\quad\tilde{\nabla}_{X}\xi=-S_{\xi}X+D_{X}\xi$
+for tangent vector fields $X$, $Y$ and normal vector field $\xi$, where $h$, $S$ and $D$ are the second fundamental form, the shape operator and the normal connection.
 The mean curvature vector field $H$ is defined by $H=(1/n){\rm trace}\,h.$ The function $|H|$ is called the mean curvature.
@@
 (2.2) $\left<h(X,Y),JZ\right>=\left<h(Y,Z),JX\right>=\left<h(Z,X),JY\right$>.
-Denote by $R$ the Riemann curvature tensor of $\nabla$. Then the equations of Gauss and Codazzi are respectively given by
+Denote by $R$ the Riemann curvature tensor of $\nabla$. Then the equations of Gauss and Codazzi are given respectively by
 $\left<R(X,Y)Z,W\right$>= $\left<h(Y,Z),h(X,W)\right>-\left<h(X,Z),h(Y,W)\right$>
@@
 $({\bar{\nabla}}_{X}h)(Y,Z)=D_{X}h(Y,Z)-h(\nabla_{X}Y,Z)-h(Y,\nabla_{X}Z).$
-3. Hamiltonian stationary Lagrangian surfaces
-Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$.
+3. Hamiltonian stationary Lagrangian surfaces with harmonic mean curvature
+Let $M$ be a Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$.
 Suppose that $H\neq 0$ everywhere.
 Denote by $K$ the Gaussian curvature of $M$.
-Let $\{e_{1},e_{2}\}$ be a local orthonormal basis of $M$ such that $Je_{1}$ is parallel to $H$.
-It follows from (2.2) that second fundamental form takes the form
+Let $\{e_{1},e_{2}\}$ be a local orthonormal frame on $M$ such that $Je_{1}$ is parallel to $H$.
+It follows from (2.2) that the second fundamental form takes the form
 (3.1) $\begin{split}&h(e_{1},e_{1})=(a-c)Je_{1}+bJe_{2},\\ &h(e_{1},e_{2})=bJe_{1}+cJe_{2},\\ &h(e_{2},e_{2})=cJe_{1}-bJe_{2}\end{split}$
-for some functions $a$, $b$ and $c$. Putting $\omega_{i}^{j}(e_{k})=\left<\nabla_{e_{k}}e_{i},e_{j}\right$>, by (2.1) and (LABEL:sf) we have
+for some functions $a$, $b$ and $c$.
+Putting $\omega_{i}^{j}(e_{k})=\left<\nabla_{e_{k}}e_{i},e_{j}\right$>, and using (2.1) and (LABEL:sf), we have
 $(\bar{\nabla}_{e_{1}}h)(e_{2},e_{2})$ $=\{e_{1}c+3b\omega_{1}^{2}(e_{1})\}Je_{1}-\{e_{1}b-3c\omega_{1}^{2}(e_{1})\}Je_{2},$
@@
 $(\bar{\nabla}_{e_{2}}h)(e_{1},e_{1})$ $=\{e_{2}(a-c)-3b\omega_{1}^{2}(e_{2})\}Je_{1}+\{e_{2}b+(a-3c)\omega_{1}^{2}(e_{2})\}Je_{2}.$
-Therefore, the equation (2.4) of Codazzi implies
+Therefore, the Codazzi equation (2.4) implies
 (3.2) $e_{1}c+3b\omega_{1}^{2}(e_{1})=e_{2}b+(a-3c)\omega^{2}_{1}(e_{2}),$
@@
 (3.5) $e_{2}a-a\omega^{2}_{1}(e_{1})=0.$
-The Hamiltonian stationary condition (1.1) is equivalent to
+Assume that $M$ is Hamiltonian stationary. Then, by (1.1) we have
 (3.6) $e_{1}a+a\omega_{1}^{2}(e_{2})=0.$
@@
 $a^{-1}e_{1},a^{-1}e_{2}]=0.$
-Therefore, there exists a local coordinate system $\{u,v\}$ such that $e_{1}=a\partial_{u}$, $e_{2}=a\partial_{v}$. Hence, the metric tensor is given by
+Therefore, there exists a local coordinate system $\{u,v\}$ such that $e_{1}=a\partial_{u}$ and $e_{2}=a\partial_{v}$. Hence, the metric tensor is given by
 (3.7) $g=a^{-2}(du^{2}+dv^{2}),$
@@
 (3.9) $K=-(a_{u})^{2}-(a_{v})^{2}+a(a_{uu}+a_{vv}).$
-Equations (3.2) and (3.3) are rewritten as
+It follows from (3.8) that (3.2) and (3.3) can be rewritten as
 (3.10) $ac_{u}+3ba_{v}$ $=ab_{v}-(a-3c)a_{u},$
 (3.11) $-ab_{u}+3ca_{v}$ $=ac_{v}-3ba_{u}.$
-Put $\delta=\epsilon-K$. Then the equation (2.3) of Gauss and (LABEL:sf) yield that
+Put $\delta=\epsilon-K$. Then the Gauss equation (2.3) together with (LABEL:sf) yields
 (3.12) $\delta=2b^{2}-ac+2c^{2}.$
+In the case where the mean curvature is constant, we have
 Proposition 3.1.
-Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$. If $|H|$ is non-zero constant, then $\bar{\nabla}h=0$.
-Proof.
-From (3.5) and (3.6), it follows that if $|H|=|a|/2$ is non-zero constant, then $\omega_{1}^{2}=0$.
-Hence we have $K=0$.
-Differentiating (3.12) along $e_{i}$ $(i=1,2)$ yields $4be_{i}b+(4c-a)e_{i}c=0$, which together with (3.2) and (3.3), shows that $b$ and $c$ are constant.
-Thus, using $\omega_{1}^{2}=0$, (2.1) and (LABEL:sf), we obtain $\bar{\nabla}h=0$.
+Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$. If $|H|$ is a non-zero constant, then $\bar{\nabla}h=0$.
+Proof. From (3.8), it follows that if $|H|=|a|/2$ is a non-zero constant, then $\omega_{1}^{2}=0$, and hence $K=0$. Equations (3.2) and (3.3) reduce to
+(3.13) $c_{u}-b_{v}=c_{v}+b_{u}=0.$
+Since $\delta=\epsilon$, differentiating (3.12) with respect to $u$ and $v$, and using (3.13), we have
+(3.14) $\begin{pmatrix}4b&4c-a\\ a-4c&4b\\ \end{pmatrix}\begin{pmatrix}b_{u}\\ c_{u}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}$.
+If $16b^{2}+(a-4c)^{2}\neq 0$ on an open subset, then by (3.14) we obtain $b_{u}=c_{u}=0$.
+Thus, it follows from (3.13) that $b$ and $c$ are constant.
+Otherwise, $b=0$ and $c$ is constant.
+In both cases, using (2.1) and (LABEL:sf), we obtain $\bar{\nabla}h=0$.
+∎
 Remark 3.1.
-An explicit description of all Lagrangian surfaces with $\bar{\nabla}h=0$ in $\tilde{M}^{2}(4\epsilon)$ has been obtained in Theorem 7.2 in [1], Theorem A.1 and Theorem A.2 in [4].
-4. Main theorem
-Let ${\mathbb{C}}^{3}_{1}$ be the complex 33-space endowed with the inner product
-$\left<(z_{1},z_{2},z_{3}),(w_{1},w_{2},w_{3})\right>={\rm Re}(-z_{1}\bar{w}_{1}+z_{2}\bar{w}_{2}+z_{3}\bar{w}_{3}).$
-Put $H_{1}^{5}(-1)=\{z\in{\mathbb{C}}^{3}_{1}:\left<z,z\right>=-1\}$.
-The map $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ defined by $z\mapsto z\cdot{\mathbb{C}}^{*}$ is referred to as the Hopf fibration.
-The main result of this paper is the following classification theorem.
+An explicit description of all Lagrangian surfaces with $\bar{\nabla}h=0$ in $\tilde{M}^{2}(4\epsilon)$ has been obtained in Theorem 7.2 of [1], Theorems A.1 and A.2 of [4].
+Our next step is to investigate the case of non-constant mean curvature. The main result of this paper is the following.
 Theorem 3.1.
-Let $M$ be a Hamiltonian stationary Lagrangian surface in a 22-dimensional complex space form $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$.
+Let $M$ be a Hamiltonian stationary Lagrangian surface in a complex space form $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$.
 Suppose that $H$ is nowhere vanishing.
-If $M$ has non-constant harmonic mean curvature and constant Gaussian curvature $K$, then $K=\epsilon=-1$ and $M$ is locally congruent to the image of $\Pi\circ\phi$, where $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ is the Hopf fibration and $\phi:M\rightarrow H_{1}^{5}(-1)\subset{\mathbb{C}}^{3}_{1}$ is given by one of the following immersions::
+If $|H|$ is a non-constant harmonic function and $K$ is constant, then $K=\epsilon=-1$ and $M$ is locally congruent to the image of $\Pi\circ\phi$, where $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ is the Hopf fibration and $\phi:M\rightarrow H_{1}^{5}(-1)\subset{\mathbb{C}}^{3}_{1}$ is given by one of the following immersions::
 $(1)$
@@
 where $m$ is a positive real number.
-Proof. We shall use the same notation as in Section 3. Suppose that $|H|$ is a harmonic function on $M$. Then by (3.7) we have
+Proof. Suppose that $|H|$ is a harmonic function on $M$. Then by (3.7) we have
 (3.15) $a_{uu}+a_{vv}=0.$
@@
 (3.16) $a_{u}=\sqrt{-K}\cos\theta,\quad a_{v}=\sqrt{-K}\sin\theta.$
-for some function $\theta$ on $M$. Substituting (4.2) into (4.1), we have
+for some function $\theta$ on $M$. Substituting (3.16) into (3.15), we get
 (3.17) $(\sin\theta)\theta_{u}+(\cos\theta)\theta_{v}=0.$
@@
 (3.18) $(\cos\theta)\theta_{u}+(\sin\theta)\theta_{v}=0.$
-It follows (4.3) from (4.4) that $\theta_{u}=\theta_{v}=0$, that is, $\theta$ is constant. Solving (4.2), we conclude that up to translations, $a$ is given by
+It follows from (3.17) and (3.18) that $\theta_{u}=\theta_{v}=0$, that is, $\theta$ is constant. Solving (3.16), we conclude that up to translations, $a$ is given by
 (3.19) $a=\sqrt{-K}\{(\cos\theta)u+(\sin\theta)v\}.$
@@
 $(a-c)c_{v}=0.$
-If $c_{v}\neq 0$ on an open subset in $\mathcal{U}$, then $a=c$ on that open subset.
+If $c_{v}\neq 0$ on an open subset in $\mathcal{U}$, then $a=c$.
 From (3.12) we see that $a$ is constant, which contradicts our assumption.
@@
 It follows from (3.20) that $a_{u}=0$ on $\mathcal{U}_{1}$.
-Hence, using (4.5) and K=1K=-1 yields that $a^{2}=v^{2}$.
-Applying the coordinate transformation: $y=-\int v^{-1}dv$, we see that the metric tensor (3.7) becomes
+Hence, using (3.19) and K=1K=-1, we obtain $a^{2}=v^{2}$.
+Applying the coordinate transformation $y=-\int v^{-1}dv$, we see that the metric tensor (3.7) becomes
 $g=m^{2}e^{2y}du^{2}+dy^{2}$
@@
 $h(\partial_{u},\partial_{u})=J\partial_{u},\quad h(\partial_{u},\partial_{y})=h(\partial_{y},\partial_{y})=0.$
-Rewriting $u$ with $x$ and according to [3, page 3475], we conclude that $\mathcal{U}_{1}$ is congruent to the Lagrangian surface obtained from (1).
+We rewrite $u$ as $x$. According to [3, page 3475], we conclude that $\mathcal{U}_{1}$ is congruent to the Lagrangian surface obtained from (1).
 Case (i.2): $a_{v}=0$ on an open subset $\mathcal{U}_{2}\subset\mathcal{U}$. In this case, $a_{u}\neq 0$. Differentiating (3.12) with respect to $u$ leads to
@@
 $(a-2c)(a-6c)a_{u}=0.$
-If $a\neq 2c$ on an open set in $\mathcal{U}_{2}$, then $a=6c$ and it follows from (3.12) that $a$ is constant, which contradicts our assumption.
+If $a\neq 2c$ on an open subset of $\mathcal{U}_{2}$, then $a=6c$, and it follows from (3.12) that $a$ is constant, which contradicts our assumption.
 Hence $a=2c$, which together with (3.12) implies $K=\epsilon=-1$.
 Thus, by (3.19) we obtain $a^{2}=u^{2}$.
-Applying the coordinate transformation: $x=(u+v)/2$ and y=(uv)/2y=(u-v)/2, we see that the metric tensor (3.7) becomes
+Applying the coordinate transformation $x=(u+v)/2$ and y=(uv)/2y=(u-v)/2, we see that the metric tensor (3.7) becomes
 $g=\dfrac{2}{(x+y)^{2}}(dx^{2}+dy^{2})$
@@
 According to [5, page 124], $\mathcal{U}_{2}$ is congruent to the Lagrangian surface obtained from (2) with $m=1$.
-Case (ii): $b\neq 0$ on an open subset $\mathcal{V}$. We put $A=a_{u}$ and $B=a_{v}$, which are constants. Differentiating (3.12) with respect to $u$ and $v$, we obtain
+Case (ii): $b\neq 0$ on an open subset $\mathcal{V}$. We put $A=a_{u}$ and $B=a_{v}$, which are constant. Differentiating (3.12) with respect to $u$ and $v$, we obtain
 (3.24) $b_{u}=\frac{cA+(a-4c)c_{u}}{4b},$
@@
 Substituting (3.24) and (3.25) into (3.10) and (3.11) yields
-(4.12) $4abc_{u}-a(a-4c)c_{v}=4b(3c-a)A-(12b^{2}-ac)B,$
-(4.13) $a(a-4c)c_{u}+4abc_{v}=(12b^{2}-ac)A+12bcB.$
-Solving (4.12) and (4.13) for $c_{u}$ and $c_{v}$ and using (3.12), we have
-(4.14) $c_{u}=\frac{Af_{1}+Bf_{2}}{a(a^{2}+8\delta)},$
-(4.15) $c_{v}=\frac{Af_{3}+Bf_{4}}{a(a^{2}+8\delta)},$
+(3.26) $a\{4bc_{u}-(a-4c)c_{v}\}=4b(3c-a)A-(12b^{2}-ac)B,$
+(3.27) $a\{(a-4c)c_{u}+4bc_{v}\}=(12b^{2}-ac)A+12bcB.$
+Solving (3.26) and (3.27) for $c_{u}$ and $c_{v}$, we have
+(3.28) $c_{u}=\frac{Af_{1}+Bf_{2}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
+(3.29) $c_{v}=\frac{Af_{3}+Bf_{4}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
 where
-$f_{1}=-4ab^{2}-a^{2}c+4ac^{2},$
-$f_{2}=-48b^{3}+16abc-48bc^{2},$
-$f_{3}=4a^{2}b+48b^{3}-32abc+48bc^{2},$
-$f_{4}=12ab^{2}-a^{2}c+4ac^{2}.$
+(3.30) $f_{1}=-4ab^{2}-a^{2}c+4ac^{2},$
+(3.31) $f_{2}=-48b^{3}+16abc-48bc^{2},$
+(3.32) $f_{3}=4a^{2}b+48b^{3}-32abc+48bc^{2},$
+(3.33) $f_{4}=12ab^{2}-a^{2}c+4ac^{2}.$
 Substituting (3.28) and (3.29) into (3.24) and (3.25) gives
-(4.16) $b_{u}=$ $\frac{Ag_{1}+Bg_{2}}{a(a^{2}+8\delta)},$
-(4.17) $b_{v}=$ $\frac{Ag_{3}+Bg_{4}}{a(a^{2}+8\delta)},$
+(3.34) $b_{u}=$ $\frac{Ag_{1}+Bg_{2}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
+(3.35) $b_{v}=$ $\frac{Ag_{3}+Bg_{4}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
 where
@@
 $g_{4}=3a^{2}b-8abc.$
-Differentiating (4.14) and (4.15) with respect to $v$ and $u$ respectively, we have
-(4.18) $c_{uv}=\frac{(Af_{1}+Bf_{2})_{v}a(a^{2}+8\delta)-(Af_{1}+Bf_{2})(3a^{2}+8\delta)B}{a^{2}(a^{2}+8\delta)^{2}},$
-(4.19) $c_{vu}=\frac{(Af_{3}+Bf_{4})_{u}a(a^{2}+8\delta)-(Af_{3}+Bf_{4})(3a^{2}+8\delta)A}{a^{2}(a^{2}+8\delta)^{2}}.$
-By a straightforward but tedious computation, using (4.14)-(4.17) and substituting (3.12) into the numerators of (4.18) and (4.19), we get
-(4.20) $c_{uv}=\frac{A^{2}h_{1}+ABh_{2}+B^{2}h_{3}}{a^{2}(a^{2}+8\delta)^{2}},$
-(4.21) $c_{vu}=\frac{A^{2}h_{4}+ABh_{5}+B^{2}h_{6}}{a^{2}(a^{2}+8\delta)^{2}},$
+Using (3.28)-(3.35), we differentiate (3.28) and (3.29) with respect to $v$ and $u$, respectively. A computation with a computer algebra system yields
+(3.36) $c_{uv}=\frac{A^{2}h_{1}+ABh_{2}+B^{2}h_{3}}{a^{2}(a^{2}+16b^{2}-8ac+16c^{2})^{2}},$
+(3.37) $c_{vu}=\frac{A^{2}h_{4}+ABh_{5}+B^{2}h_{6}}{a^{2}(a^{2}+16b^{2}-8ac+16c^{2})^{2}},$
 where
@@
 $h_{6}=$ $-240a^{2}b^{3}+80a^{3}bc+768ab^{3}c-496a^{2}bc^{2}+768abc^{3}.$
-Since $c_{uv}-c_{vu}$=0 holds, from (4.20), (4.21), it follows that
-(4.22) $4b(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})=0,$
+Subtracting (3.37) from (3.36), we have
+(3.38) $c_{uv}-c_{vu}=\frac{4b(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})}{a^{2}(a^{2}+16b^{2}-8ac+16c^{2})^{2}},$
 where
@@
 $+132a^{2}c^{2}+384b^{2}c^{2}-288ac^{3}+192c^{4}.$
-By a straightforward but tedious computation, using (4.14)-(4.17), we get
-(4.23) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{u}$ $=\frac{4(A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4})}{a^{2}+8\delta},$
-(4.24) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{v}$ $=\frac{8(A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8})}{a^{2}+8\delta},$
+[…]
+(3.39) $A^{2}k_{1}+ABk_{2}+B^{2}k_{3}=0.$
+Differentiating the left hand side of (3.39) and using (3.28)-(3.35), with the aid of a computer algebra system, we get
+(3.40) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{u}$ $=\frac{4(A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4})}{a^{2}+16b^{2}-8ac+16c^{2}},$
+(3.41) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{v}$ $=\frac{8(A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8})}{a^{2}+16b^{2}-8ac+16c^{2}},$
 where
@@
 (3.43) $A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8}=0.$
-Using a computer algebra system, we check that the resultant of the left-hand sides of (4.25) and (4.26) with respect to AA is, up to a constant factor,
+Using a computer algebra system, we find that the resultant of the left-hand sides of (3.42) and (3.43) with respect to AA is, up to a non-zero constant factor,
 (3.44) $B^{9}(2b^{2}-ac+2c^{2})^{3}(a^{2}+16b^{2}-8ac+16c^{2})^{7}(a^{2}+48b^{2}-24ac+48c^{2})^{2}Q,$
@@
 $-6912ab^{2}c^{3}+2448a^{2}c^{4}+6912b^{2}c^{4}-3456ac^{5}+2304c^{6}.$
-Combining (3.12) and (4.27), we find that (4.27) can be simplified to
-(4.28) $B^{9}\delta^{3}(a^{2}+8\delta)^{7}(a^{2}+24\delta)^{2}(5a^{6}+60a^{4}\delta+180a^{2}\delta^{2}+288\delta^{3}),$
-which must vanish on $\mathcal{V}$. Taking into account $\nabla a\neq 0$, we have
+Combining (3.12) and (3.44), we find that (3.44) simplifies to
+$B^{9}\delta^{3}(a^{2}+8\delta)^{7}(a^{2}+24\delta)^{2}(5a^{6}+60a^{4}\delta+180a^{2}\delta^{2}+288\delta^{3}),$
+which must vanish on $\mathcal{V}$. Taking into account that $\nabla a\neq 0$, we have
 $B\delta=0.$
@@
 which contradicts $\nabla a\neq 0$. Hence $\delta=0$, that is, $K=\epsilon=-1$.
-From (3.12) we see that $c\neq 0$. Changing the sign of $e_{1}$ if necessary, we may assume that $c>0$. Put
+From (3.12) and $b\neq 0$, we see that $c\neq 0$. Changing the sign of $e_{1}$ if necessary, we may assume that $c>0$. Put
 (3.48) $f=\frac{\sqrt{b^{2}+c^{2}-b\sqrt{b^{2}+c^{2}}}}{\sqrt{2}(b^{2}+c^{2})},\quad k=\frac{\sqrt{b^{2}+c^{2}+b\sqrt{b^{2}+c^{2}}}}{\sqrt{2}(b^{2}+c^{2})},$
@@
 (3.49) $a=\frac{\sqrt{f^{2}+k^{2}}}{fk},\quad b=\frac{k^{2}-f^{2}}{(f^{2}+k^{2})^{\frac{3}{2}}},\quad c=\frac{2fk}{(f^{2}+k^{2})^{\frac{3}{2}}}.$
-We make the following change of the basis:
+We make the following change of basis:
 (3.50) $\tilde{e}_{1}=\frac{k}{\sqrt{f^{2}+k^{2}}}e_{1}+\frac{f}{\sqrt{f^{2}+k^{2}}}e_{2},\quad\tilde{e}_{2}=\frac{f}{\sqrt{f^{2}+k^{2}}}e_{1}-\frac{k}{\sqrt{f^{2}+k^{2}}}e_{2}.$
@@
 $h(\tilde{e}_{1},\tilde{e}_{1})=f^{-1}J\tilde{e}_{1},\quad h(\tilde{e}_{1},\tilde{e}_{2})=0,\quad h(\tilde{e}_{2},\tilde{e}_{2})=k^{-1}J\tilde{e}_{2}.$
-Therefore, according to [5, Theorem 4.1], there exists a local coordinate system $\{x,y\}$ such that $\partial_{x}=f\tilde{e}_{1}$ and $\partial_{y}=k\tilde{e}_{2}$, where $f$ and $k$ satisfy
+According to [5, Theorem 4.1], there exists a local coordinate system $\{x,y\}$ such that $\partial_{x}=f\tilde{e}_{1}$ and $\partial_{y}=k\tilde{e}_{2}$, where $f$ and $k$ satisfy
 (3.51) $\frac{f_{y}}{k}=\frac{k_{x}}{f},\quad\Bigl(\frac{f_{y}}{k}\Bigr)_{y}+\Bigl(\frac{k_{x}}{f}\Bigr)_{x}=fk.$
@@
 (3.55) $f=\frac{m\sqrt{1+m^{2}}}{m^{2}x+y},\quad$ $k=\frac{\sqrt{1+m^{2}}}{m^{2}x+y},$
-where $\lambda$ and $m$ are positive real numbers.
-We note that $f=mk$. By the first equation of (4.33) we have
+where $\lambda$ and $m$ are positive real numbers. We note that $f=mk$. By the first equation of (3.49) we have
 $a=\frac{\sqrt{1+m^{2}}}{mk}.$
@@
 In [6, Section 7], it is proved that the corresponding surface is congruent to the Lagrangian surface obtained from case (2) with $m\neq 1$.
-This completes the proof.
+∎